2. Use established properties of moduli to show that when |23| + |24|| Re(21 + z2) |z3 + z4| |z1| + |22|| - ||z3| – |z4|||
Answer #1

we proove That Re(21422), 12, 1+122) 123 tzal 11231-12411 We know theit Re(2) < 121 1231 € 1241 . Recai) < 12,1 & Re (22) < 1Noco consider lza1 = 12a + 2z - 231 = 123+24+ (-23) 1. (-23) 1 = 123tzult 1231 (( By triangle inequality 121+22121211+1221 9/